LINE L1-Y-X=0,;L2=2X+Y INTERSECT L3.;Y+2=0 AT P AND Q THE BISECTOR OF ACUTE ANGLE BETWEEN L1 AND L2 INTERSECT L3 AT R.. IS THE FDOLLOWING STATEMENT TRUE OR NOT IF YES WHY?? THE RATIO OF PR:RQ IS 2ROOT2:ROOT5??

LINE ย L1-Y-X=0,;L2=2X+Y INTERSECT L3.;Y+2=0 ย AT P AND Q


THE BISECTOR OF ย ACUTE ANGLE BETWEEN L1 AND L2 INTERSECTย 


L3 AT R.. IS THE FDOLLOWING STATEMENT TRUE OR NOT IF YES WHY??


THE RATIO OF PR:RQ IS 2ROOT2:ROOT5??

Grade:11

1 Answers

shaswat roy
13 years ago

eqn of d angle bisector will b (|y-x|)/root2=(|y+2x|)/root5

logically it can b seen d eqn of d acute angle bisector will be: (y-x)/root2=(y+2x)/root5  (1)                   [removin mod sign]

thus d pt R can b obtained by solving(1) and y=-2 which will give [(-2)(root5-root2)/2root2+root5),-2]

P(-2,-2), Q(1,-2)

now find out PR nd RQ as we know deir coordinates

on solvingPR:RQ comes out to b 2root2:root5,hence d statement is true